Percent of Sales Method

At the end of the current year, Accounts Receivable has a balance of $465,000; Allowance for Doubtful Accounts has a debit balance of $4,000; and sales for the year total $2,090,000. Bad debt expense is estimated at 1/4 of 1% of sales

a. Determine the amount of the adjusting entry for uncollectible accounts

b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Bad Debt Expense

Accounts Receivable $
Allowance for Doubtful Accounts $
Bad Debt Expense

c. Determine the net realizable value of accounts receivable
Answer #1

a) Amount of the adjusting entry = 2090000*1%*1/4 = 5225

b) Account receivable = 465000

Allowance for doubtful accounts = 5225-4000 = 1225

Bad debt expense = 5225

c) Net realizable value = 465000-1225 = 463775